A car moving along a straight highway with speed of 126 km h -1 is
brought to a stop within a distance of 200 m. What is the retardation of
the car (assumed uniform), and how long does it take for the car to
stop?
(a) -3.06 m/s 2 ,11.44s
(b) 3.06 m/s 2 ,1.44s
(c) 36 m/s 2 ,1.44s
(d) 42.06 m/s 2 ,31.44s

Grade:11

1 Answers

Vikas TU
7 years ago
From third law of eqn. of motion we get,
v^2 = u^2 + 2as
0^2 = (126000/3600)^2 + 2a*200
we get
a = –(1260/36)^2/400
   = – 3.0625 m/s^2.
 
and time=>
0 = 1260/36 – 3.0625*t
t = 11.44 seconds